Solve using the five-step problem-solving process. Show all steps necessary to arrive at your solution.
You are putting a stone border of uniform width around a rectangular garden that measures 6 yards by 15 yards. You only have enough stone to cover 100 square yards. How wide should the border be?

can't lie I don't know what kinda of problem this is: can you explain and help? Word problems confuse me everytime.

Answer by mananth(16949)

Let x be the width of the stone border.
Including the stone border trhe area will be
(6+2x)(15+2x)
The area of garden = 15*6=90 sq. yards
..
Arae including width minus area of garden = area of stone border.
(6+2x)(15+2x)-90 =100
90+12x+30x+4x^2- 90=100
4x^2+42x-100=0
4x^2+50x-8x-100=0
4x(x+12.5)-8(x+12.5)=0
(4x-8)((x+12.5)=0
4x=8
or x=2
..
x=-12.5 .. ignore this value